OpenStudy (anonymous):

If r>0 and r^t = 6.25r^(t+2), then r = a) 2/5 b) 4/9 c) 5/8 d) 3/4 e) 5/4

OpenStudy (whpalmer4):

Answer will be a fraction, so write 6.25 as a fraction: 25/4 \[r^t = \frac{25}{4}r^{t+2}\]Divide both sides by \(r^t\) \[1 = \frac{25}{4}r^2\]Take square root of both sides\[1 = \sqrt{\frac{25}{4}}*\sqrt{r^2}\]Simplify and solve for \(r\)